OpenStudy (anonymous):

the product of two consecutive odd integers is equal to 27 plus six times the sum of the integers. What is the larger integer?

OpenStudy (anonymous):

\[n(n+2)=27+6(n+n+2)\] is more like it

OpenStudy (anonymous):

once I factor everything out use quadratic formula?

OpenStudy (anonymous):

\[n^2+2n=27+6(2n+2)\] \[n^2+2n=27+12n+12\] \[n^2-10n-39=0\]

OpenStudy (anonymous):

you can factor this one

OpenStudy (anonymous):

\[(n-13)(n+3)=0\] etc

OpenStudy (anonymous):

13 or -3